OpenStudy (anonymous):

how do you put 27 miles and 1/2 an hour in together as a unit rate

OpenStudy (anonymous):

Well you rate is also speed so, \[\frac{ distance }{ time }=\frac{ miles }{ minutes }=\frac{ 27 }{ 30 }=\frac{ 9 }{ 10 }miles/minute\]

OpenStudy (anonymous):

You can put it into hours as well

OpenStudy (anonymous):

opk but the only options are 13.5 miles per hour 44 miles per hour 54 miles per hour and 56 miles per hour

OpenStudy (anonymous):

So if you put it in hours, it'll be\[\frac{ distance }{ time }=\frac{ miles }{ hours }=\frac{ 27 }{ .5 }=13.5 miles/hour\]

OpenStudy (anonymous):

Oh wait my bad its 54 miles per hour, I accidently times .5 instead of dividing.

OpenStudy (anonymous):

oh ok thank you